if you used nichrome wire .4m in length, and .02 cm in diameter, to make a small solenoid. and if you used a 1.4 v battery, what would be the direction and magnitude of the magnetic field at the centre of this solenoid?
1
Expert's answer
2018-11-21T15:14:09-0500

ρ=110×〖10〗^(-8) Ohm∙m

l=4 m

d=2×〖10〗^(-4) m

U=1.4 V

Current I= U/R , where R= ρ l/S, then I=US/ρl , S= π d^2/4, I=(πUd^2)/4ρl

I=(1.4 ∙3.14 ∙ (2×〖10〗^(-4) )^2)/(4∙110×〖10〗^(-8)∙4)=0.01(A)

If the radius of the solenoid is r = 1 cm=0.01 m, then B=(μ_0 I)/2πr=(4π×〖10〗^(-7)∙0.01)/(2π∙0.01)=2×〖10〗^(-7) Wb